find the length of the hypotenuse, to the nearest tenth of a centimeter, of a right triangle if one angle measures 70 degrees and the adjacent leg measures 8 cm?
Answer by Alan3354(69443)

cosine = adj/hyp
cos(70) = 8/hyp
hyp = 8/cos(70)
hyp =~ 23.4 cm